You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100217 (19 Feb 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 February16-17

Observer: V. Testa, K.Boutsia, F. Pedichini
Telescope Operator: D.Gonzales Huerta
Support Astronomer: D. Thompson
Instrument: LUCIFER 1

Summary

2.5h lost for telescope+AGw problems
Night of fantastic quality, both great seeing and probably photometric except in the last hour or so
Mostly preimaging for MOS masks preparation

Data Summary

Project Objects Filter ExpTime Notes
Turatto SN2009kn H, Ks 20 + 20min  
Zamorani Cosmos field Ks 15' x 2 preimaging for MOS
Grazian GOODS-N Ks 14' x 4 preimaging for MOS
Mannucci HDF-N Ks 14' preimaging for MOS
Mannucci Q1422 J, Ks 60' + 105' preimaging + science

Details

00:45 - Start Br-gamma sky flats
  • ~8500 cts; image Nos. 0012-0021
01:12 - Low Br-gamma flats
  • ~2000 cts 0022-0031

01:15 - High Ks flats *~9000 cts 0032-0041
01:22 - Low Ks flats
  • ~2000cts 0042-0051

01:26 - High J flats
  • ~10000 cts 0052-0061
01:28 - Low J flats
  • ~2000 cts 0062-0071

01:33 - Correct pointing
  • High elev. WT10_314 0072-0073
  • Low elev. WT10_078 0074-0075 new IA=-370
  • Go back to WT10_314 0076-0077

01:47 - Go to Persson standard to collimate
  • BS9108 (POSITION,ACTIVE) problems in presetting AGw has problems

Lost one hour for problems with AGw

02:50 - start again, strongly out of focus, doing focus by hand to reduce and put into reach of guide probe

Trying to do spectroscopic target. Telescope problems. Lost in total 2.5h.

04:05 - Persson standard (BS9138) close to target 0094-0101

04:25 - Go to Cosmos field for preimaging.

04:27 - field CL4: 6x10s 15 jitters 20" radius 0102-0117

05:17 - field CLA1: 6x10s x 10 jitters 20" radius 0119-0128

05:41 - Guide star out of the field. Correcting pointing
  • WT10_220 0129-0130

05:49 - field CLA2: 6x15s x 10 jitters 20" radius 0132-0143

06:17 - Turatto SN2009kn Ks imaging 20 min: 6x15sx10 jitters 40" radius 0145-0154

06:58 - Turatto SN2009kn H imaging 20 min: 6x15sx10 jitters 40" radius 0157-0236 (8x15x10-normal????)

07:29 - GOODS-N1 (15 minutes lost for collimation): 6x20secx6 jitters 20" radius 0238-0243

08:07 - slew to GOODS-N2 -collimation + pointing correction

08:15 - GOODS-N2 : 6x20sx7 jitters 20" radius 0247-0253

08:40 - GOODS-N4 : 6x20sx7 jitters 20" radius 0255-0261

09:27 - GOODS-N3 : 6x20sx7 jitters 20" radius 0263-0269

10:00 - HDF - N: 6x20sx7 jitters 20" radius 0271-0277

10:30 - Q1422 - Ks: 6x15sx40 jitter 40 radius 0280 - 338

11:08 - stopped because sources were elongated. Pointing corrected. Exposure resumed
  • Q1422 - Ks: 6x15sx30 jitter 40 radius 0342 - 0371 (we stopped when 30/40 exposures were done)

12:14 - Q1422 - J: 3x30sx30 jitter 30 radius 0372 - 0401

13:14 - Q1422 - Ks: 6x15sx30 jitter 40 radius = 0402 - 405=

13:45 Dark Images
  • 10x6x10
  • 15x6x10
  • 20x6x10
  • 30x3x10

14:00 End of Night

-- VincenzoTest - 17 Feb 2010
Topic revision: r5 - 19 Feb 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback